OpenStudy (anonymous):

Solve this system algebraically. 9x + 2y = 5 y - 2x + 3 = 0 {(11/13, -17/13)} {(-11/13, -17/13)} {(-11/13, 17/13)} {(11/13, 17/13)}

OpenStudy (anonymous):

9x + 2(2x-3) = 5 13x = 11 x = 11/13

OpenStudy (anonymous):

y = 22/13 - 3 = -17/13

OpenStudy (anonymous):

What is the value of the system determinant for the following system of equations? 6x + y = 4 5x + y - 7 = 0 -1 1 31

OpenStudy (anonymous):

1

OpenStudy (anonymous):

x + y = 12 x - y = 10 What is the value of the x-determinant for the system shown? -22 -12 -2